# Deep Operator Networks for Bayesian Parameter Estimation in PDEs

### v1

@@ Abstract
We present a novel framework combining Deep Operator Networks (DeepONets) with Physics-Informed Neural Networks (PINNs) to solve partial differential equations (PDEs) and estimate their unknown parameters. By integrating data-driven learning with physical constraints, our method achieves robust and accurate solutions across diverse scenarios. Bayesian training is implemented through variational inference, allowing for comprehensive uncertainty quantification for both aleatoric and epistemic uncertainties. This ensures reliable predictions and parameter estimates even in noisy conditions or when some of the physical equations governing the problem are missing. The framework demonstrates its efficacy in solving forward and inverse problems, including the 1D unsteady heat equation and 2D reaction-diffusion equations, as well as regression tasks with sparse, noisy observations. This approach provides a computationally efficient and generalizable method for addressing uncertainty quantification in PDE surrogate modeling.
@@
journal: arXiv
@@
[csulb]organization=Depratment of Computer Engineering and Computer Science, addressline= California State University, city=Long Beach, postcode=90840, state=CA, country=USA
@@ 1 Introduction
Physics-Informed Neural Networks (PINNs) [7, 8] integrate governing physical laws, expressed as PDEs, directly into their loss functions. This approach enables PINNs to solve forward and inverse problems efficiently, even with limited data, by ensuring that predictions adhere to the underlying physics. By embedding physical laws into the learning process, PINNs overcome the reliance on extensive datasets required by traditional neural networks, making them more practical for real-world scenarios [9]. Applications of PINNs include solving fluid dynamics problems [10, 11], addressing inverse problems [12, 13], controlling dynamical systems [14, 15, 16], and performing uncertainty quantification [17].
@@
In scientific machine learning, probabilistic models are highly valued for estimating noise in observational data, addressing model errors, and accounting for incomplete knowledge of the underlying physics. Viana et al. [23] highlight the evolution of scientific modeling, emphasizing the role of physics-informed neural networks in reducing computational costs and enhancing modeling flexibility. Molnar et al. [24] demonstrate the application of PINNs within a Bayesian framework to reconstruct flow fields from projection data, showcasing their ability to incorporate physics priors and improve accuracy in noisy environments. Li et al. [25] explore the challenges of solving inverse problems with PINNs, advocating for a Bayesian approach to provide robust uncertainty quantification, particularly when observations are sparse or noisy.
@@
Bayesian Neural Networks (BNNs) [26, 27, 28] provide a probabilistic framework for uncertainty quantification by assigning distributions to network weights. While posterior estimation over BNN parameters is challenging due to high-dimensionality and non-convexity, methods such as Hamiltonian Monte Carlo (HMC) and variational inference (VI) approximate the posterior effectively, albeit at higher computational costs than maximum likelihood approaches. Yang et al. [29] introduced Bayesian Physics-Informed Neural Networks (B-PINNs), combining BNNs and PINNs to address forward and inverse PDE problems in noisy conditions. By integrating physical equations into training, B-PINNs offer a practical solution for probabilistic PDE learning tasks.
@@
Accurately estimating PDE parameters while quantifying uncertainty remains a significant challenge, particularly in complex systems with scarce or noisy data. The high computational demands and challenges in selecting suitable priors in Bayesian approaches restrict the practical use of black-box sampling methods in scientific machine learning. This presents an opportunity for new methodologies that effectively combine physics-guided training with computationally efficient Bayesian approaches.
@@
In this work, we propose a novel method that extends the capabilities of PINNs by integrating Bayesian methodologies with Deep Operator Networks (DeepONets) to enhance parameter estimation in PDEs. Our approach relies on variational inference: starting with a prior distribution for the unknown PDE parameters, we apply Bayesian inference to estimate the posterior distributions based on available data and physics. The DeepOperator architecture modulates the forward prediction of the solution of the PDE with inverse parameter estimation. This method offers substantial computational advantages, providing uncertainty quantification comparable to B-PINNs but with significantly reduced computational costs.
@@
We continue our discussion in the following steps: In section 2, we detail the methodology of Bayesian Deep Operator Networks, demonstrate their application to parameter estimation in PDEs, and discuss their advantages and shortcomings compared to existing approaches. Section 3 presents experimental results on a variety of PDEs, showcasing the effectiveness of our method in solving forward and inverse problems. Finally, we discuss the implications of our findings for future research and potential applications in various scientific and engineering domains in section 4.
@@ 2.1 Problem definition
We start by considering a parametrized partial differential equation (PDE), defined over time span $t\in[t_{0},t_{f}]$ and spatial domain $x\in\Omega$:
@@
we are interested in training a probabilistic machine learning model that simultaneously approximates $y(t,x)$ and $\lambda$ based on the data eq. 2 and the PDE eq. 1. We will start from a Bayesian formulation of this problem and derive a corresponding loss function for training the model.
The posterior distribution over the latent parameter $\lambda$ given data $\mathcal{D}$ is expressed as
@@
The likelihood term $p(\mathcal{D}\mid\lambda)$ is constructed from two components: the data likelihood $p(y_{i}\mid t_{i},x_{i},\lambda)$, which captures the relationship between the inputs $t_{i},x_{i}$ and outputs $y_{i}$, and the likelihood of the predicted solution satisfying the governing equations. The latter is achieved via the residuals of eq. 1.
@@
where $\sigma_{R}^{2}$ is the variance of the residual noise. $\mathcal{R}_{\text{PDE}}$ can either be simplified as $|F(t_{i},x_{i},y_{i},\lambda)|$ or modified to include other constraints such as the boundary and initial conditions. The corresponding log-likelihood of the residuals is
@@
We sample the posterior $p(\lambda\mid\mathcal{D})$ in equation eq. 3 using variational inference (VI). Considering an approximate posterior distribution $q(\lambda)$ we minimize the Kullback-Leibler (KL) divergence:
@@
Substituting eq. 3 for $p(\lambda\mid\mathcal{D})$ and rearranging terms, we obtain the well-known Evidence Lower Bound (ELBO) objective function:
@@
Monte Carlo approximation of the integral in eq. 10 yields Expanding $p(\mathcal{D}\mid\lambda)$ using the joint likelihood, eq. 10 becomes
@@
Equation 12 can be evaluated using Monte Carlo sampling of the latent variable $\lambda$ from the posterior $q(\lambda)$. The observation and residual noise variances $\sigma_{y}^{2}$ and $\sigma_{R}^{2}$ can be considered constants or learned during training. This enables a flexible uncertainty quantification method that can be utilized to estimate model or data uncertainty. We experiment with different variance estimation strategies in the experiments section 3.
The KL divergence term in eq. 12 can be approximated empirically or analytically: If both the prior $p_{0}(\lambda)$ and the variational distribution $q(\lambda)$ are normally distributed, with means $(\mu_{0}=0,\mu_{q})$ and variances $(\sigma_{0}^{2}=1,\sigma_{q}^{2})$, the KL divergence can be written as:
@@ 2.2 Model architecture
Figure 1: DeepBayONet architecture.
$\mathcal{M}_{\theta}(t,x;\lambda)$ is a deep operator network (DeepONet) that consists of two components: the branch network and the trunk network [30, 31]. The branch network approximates the forward solution $y(t,x)$ of the PDE. The trunk network processes independent prior samples to generate samples of the approximate posterior $q(\lambda)$ that also influence the branch network’s predictions. See fig. 1 for a schematic representation of the network architecture. Our proposed architectue, hereafter referred to as DeepBayONet uses the DeepONet structure to estimate the forward solution and parameters of the PDE at the same time:
@@
where $W_{0}$ and $b_{0}$ are the output layer weights and bias, $\sigma(\cdot)$ is the activation function, and $\odot$ denotes the element-wise product. The branch network $\mathsf{b}(t,x)$ and trunk network $\mathsf{\tau}(\lambda)$ are general parametrized functions that can take the form of any neural network architectures including convolutional, recurrent, or transformer networks. The entire set of trainable parameters in the model is denoted by $\theta$.
@@
Uncertainty estimation is critical in deep neural networks, particularly in scientific modeling and prediction tasks, to assess the reliability of predictions and the robustness of the model. There are two sources of uncertainty that can be considered: Aleatoric uncertainty arises from inherent noise in the observations, reflecting variability that cannot be reduced even with additional data. This type of uncertainty is often captured by modeling the variance of the output distribution, as demonstrated in probabilistic models such as heteroscedastic neural networks [32]. In contrast, epistemic uncertainty reflects the model’s lack of knowledge or suffiecient training about the underlying data distribution and can be reduced by incorporating additional data or improving the model architecture. Bayesian neural networks (BNNs) and dropout are among various methods that cab address epistemic uncertainty [32] by learning a distribution over the model parameters, allowing the model to express uncertainty in regions of the input space the model’s predictions are ambiguous. These two sources of uncertainty provide complementary information: aleatoric uncertainty quantifies noise in the data, while epistemic uncertainty assesses the model’s confidence in its predictions.
@@
In the proposed model, both aleatoric and epistemic uncertainties are quantified explicitly through different components of the loss function and the model’s probabilistic formulation. Aleatoric uncertainty is captured through the observation noise variance $\sigma_{y}^{2}$, which governs the data likelihood term in the evidence lower bound (ELBO) formulation eq. 12 where $\sigma_{y}^{2}$ represents the variance of the observation noise. This term accounts for the inherent variability in the measurements, and by treating $\sigma_{y}^{2}$ as a learnable parameter, the model can adaptively estimate the aleatoric uncertainty directly from the data. High values of $\sigma_{y}^{2}$ indicate regions with significant observation noise, leading to reduced confidence in the predictions.
Epistemic uncertainty, on the other hand, is inferred from the posterior distribution $q(\lambda)$ over the latent parameters $\lambda$, as derived through variational inference. The posterior distribution is approximated by minimizing the Kullback-Leibler (KL) divergence in eq. 12. The variability in $q(\lambda)$, often characterized by its variance $\sigma_{q}^{2}$, quantifies the model’s uncertainty about $\lambda$. Furthermore, the residual noise variance $\sigma_{R}^{2}$, associated with the residuals of the governing equation eq. 8 also contributes to epistemic uncertainty by capturing how well the predicted solutions satisfy the physical constraints imposed by the partial differential equations (PDEs).
Through Monte Carlo sampling of eq. 12, the model evaluates both aleatoric and epistemic uncertainties simultaneously, enabling a robust and interpretable quantification of uncertainty in predictions.
2.3 Relation to Dropout
A distinctive feature of the DeepBayONet architecture is the interaction between the branch and trunk networks, which not only enables the model to represent the forward solution but also plays a role in uncertainty quantification. Specifically, the branch network $\mathsf{b}(t,x)$ approximates the forward operator by mapping the inputs $t$ and $x$ to a latent feature space. Simultaneously, the trunk network $\mathsf{\tau}(\lambda)$ processes samples of the latent parameter $\lambda$ drawn from the approximate posterior distribution $q(\lambda)$. These two networks interact through an element-wise product in eq. 14. The stochastic sampling of $\lambda\sim q(\lambda)$ introduces variability in $\mathsf{\tau}(\lambda)$, thereby perturbing the forward operator represented by $\mathsf{b}(t,x)$ during training.
This mechanism of perturbation bears a conceptual resemblance to dropout [33], widely-used regularization as well as epistemic (model) uncertainty quantification technique in neural networks. Dropout introduces noise during training by randomly setting activations to zero with a probability sampled from a Bernoulli distribution [33]. In DeepBayONet, instead of discrete binary perturbations, the trunk network introduces continuous perturbations drawn from the posterior distribution $q(\lambda)$. This continuous modulation of the branch network activations improves robustness while allowing the model to incorporate epistemic uncertainty in its predictions. Sampling from $q(\lambda)$ enables the model to account for uncertainty in the latent parameters $\lambda$. This uncertainty is propagated through the interaction between the branch and trunk networks. This continuous perturbation mechanism has two significant implications. First, it ensures that the model’s predictions generalize well by preventing the branch network from overfitting to deterministic solutions. Second, it provides a means to quantify epistemic uncertainty explicitly, as the variability in $\mathsf{\tau}(\lambda)$ represents the model’s confidence in its parameter estimates. Thus, the DeepONet architecture naturally integrates regularization and uncertainty quantification in a unified framework, leveraging the interaction between the branch and trunk networks to achieve both objectives.
@@ 2.4 Practical Aspects of the Loss Function
The loss function defined in eq. 12 enforces adherence to the governing equations of the dynamical system while capturing uncertainties in both the observations and the parameters. To ensure practical implementation, the total loss is decomposed into distinct components, each addressing specific aspects of the problem. Assuming $\mathcal{M}_{\theta}(t_{i},x_{i};\lambda)=\tilde{y}_{i}$ is the predicted PDE solution, the total loss is expressed as:
$\mathcal{L}=w_{\text{Interior}}L_{\text{Interior}}+w_{\text{IC}}L_{\text{IC}}+w_{\text{BC}}L_{\text{BC}}+w_{\text{STD}}L_{\text{STD}},$
where the individual loss components are defined to correspond with the probabilistic formulation in eq. 12.
To align these loss terms with the evidence lower bound (ELBO), variance-based scaling is introduced to reflect the uncertainties in the observations and residuals. The revised definitions of the individual components are:
@@
$L_{\text{STD}}(\theta,\lambda)$ $=\log{\sigma_{R}^{2}}+\log{\sigma_{y}^{2}}$ ) (15d)
where $\sigma_{R}^{2}$ is the variance of the residual noise and $\sigma_{y}^{2}$ is the variance of the observation noise. These variances play a critical role in balancing the contributions of the individual loss terms, ensuring that the model accounts for uncertainties appropriately.
The weights $w_{\text{Interior}},w_{\text{IC}},w_{\text{BC}},$ and $w_{\text{STD}}$ are hyperparameters used to balance the influence of the respective loss components. These weights can be selected through hyperparameter optimization or by employing adaptive weighting techniques, as explored in [34, 35]. The inclusion of variance scaling in eqs. 15a, 15b and 15c ensures consistency with the ELBO formulation, connecting the practical implementation to the probabilistic foundations of the model. This alignment also allows the model to accurately capture both aleatoric and epistemic uncertainties during training and inference.

### v2

@@ Abstract
We present a novel framework combining Deep Operator Networks (DeepONets) with Physics-Informed Neural Networks (PINNs) to solve partial differential equations (PDEs) and estimate their unknown parameters. By integrating data-driven learning with physical constraints, our method achieves robust and accurate solutions across diverse scenarios. Bayesian training is implemented through variational inference, allowing for comprehensive uncertainty quantification for both data and model uncertainties. This ensures reliable predictions and parameter estimates even in noisy conditions or when some of the physical equations governing the problem are missing. The framework demonstrates its efficacy in solving forward and inverse problems, including the 1D unsteady heat equation, 2D reaction-diffusion equations, 3D eigenvalue problem, and various regression tasks with sparse, noisy observations. This approach provides a computationally efficient and generalizable method for addressing uncertainty quantification in PDE surrogate modeling.
@@
journal: Computer Physics Communications
@@
[csulb]organization=Department of Computer Engineering and Computer Science, addressline= California State University, city=Long Beach, postcode=90840, state=CA, country=USA
@@ 1 Introduction
Physics-Informed Neural Networks (PINNs) [7, 8] integrate governing physical laws, expressed as PDEs, directly into their loss functions. This approach enables PINNs to solve forward and inverse problems efficiently, even with limited data, by ensuring that predictions adhere to the underlying physics. By embedding physical laws into the learning process, PINNs overcome the reliance on extensive datasets required by traditional neural networks, making them more practical for real-world scenarios [9]. Applications of PINNs include solving fluid dynamics problems [10, 11], inverse problems [12, 13], controlling dynamical systems [14, 15, 16], and performing uncertainty quantification [17].
@@
In scientific machine learning, probabilistic models are highly valued for estimating noise in observational data, handling model errors, and accounting for incomplete knowledge of the underlying physics. Viana et al. [23] highlight the evolution of scientific modeling, emphasizing the role of physics-informed neural networks in reducing computational costs and enhancing modeling flexibility. Molnar et al. [24] demonstrate the application of PINNs within a Bayesian framework to reconstruct flow fields from projection data, showcasing their ability to incorporate physics priors and improve accuracy in noisy environments. Li et al. [25] explore the challenges of solving inverse problems with PINNs, advocating a Bayesian approach to provide robust uncertainty quantification, particularly when observations are sparse or noisy.
@@
Bayesian Neural Networks (BNNs) [26, 27, 28] provide a probabilistic framework for uncertainty quantification by assigning distributions to network weights. While posterior estimation over BNN parameters is challenging due to high-dimensionality and non-convexity, methods such as Hamiltonian Monte Carlo (HMC) and variational inference (VI) approximate the posterior effectively, albeit at higher computational costs than maximum likelihood approaches. Yang et al. [29] introduced Bayesian Physics-Informed Neural Networks (B-PINNs), combining BNNs and PINNs to address forward and inverse PDE problems in noisy conditions. By integrating physical equations into training, B-PINNs offer a practical solution to probabilistic PDE learning tasks.
Accurately estimating PDE parameters while quantifying uncertainty remains a significant challenge, particularly in complex systems with scarce or noisy data. While Deep Operator Networks [30, 5, 8] can learn PDE solution operators from data, they provide deterministic predictions and lack a principled mechanism for uncertainty quantification. To resolve this limitation, B-DeepONet [31] extends the DeepONet framework by introducing a Bayesian formulation, enabling posterior sampling over the solution operator space. B-DeepONet learns model uncertainty by learning a posterior distribution over model parameters using Langevin dynamics. During training, a variant of Stochastic Gradient Langevin Dynamics (SGLD) applies gradients alternatively to the parameters of the branch and trunk networks to help reduce the computational cost of calculating gradients for the ensemble of parameters.
@@
The high computational demands and challenges in selecting suitable priors in Bayesian approaches restrict the practical use of black-box sampling methods in scientific machine learning. This presents an opportunity for new methodologies that effectively combine physics-guided training with computationally efficient Bayesian approaches.
@@
In this work, we propose a novel method that extends the capabilities of PINNs by integrating Bayesian methodologies with Deep Operator Networks (DeepONets) [32] to improve parameter estimation in PDEs. Our approach relies on variational inference: starting with a prior distribution for the unknown PDE parameters, we apply inference to estimate the posterior distribution of the PDE solution and the parameters based on the available data and known physics of the problem. The DeepOperator architecture modulates the forward prediction of the solution of the PDE with inverse parameter estimation. This method offers flexible uncertainty quantification and accurate differentiable solutions.
@@
We continue our discussion in the following steps: In section 2, we detail the methodology of Bayesian Deep Operator Networks, demonstrate their application to parameter estimation in PDEs, and discuss practical ways of quantifying model and data uncertainty in their predictions. Section 3 presents experimental results on a variety of PDEs, which showcase the effectiveness of our method in solving forward and inverse problems. Finally, we discuss the implications of our findings for future research and potential applications in various scientific and engineering domains in section 4. The source code for these experiments is publicly available as Jupyter Notebooks at https://github.com/csml-beach/deep-bayesian-operator-nets.
@@ 2.1 Problem definition
We start by considering a partial differential equation (PDE) with a latent parameter $\lambda$ defined over the time span $t\in[t_{0},t_{f}]$ and spatial domain $x\in\Omega$:
@@
we are interested in training a probabilistic machine learning model that simultaneously approximates $y(t,x)$ and $\lambda$ based on the data eq. 2 and the PDE eq. 1. We will start from a Bayesian formulation of this problem and derive a corresponding variational loss function for training the model.
The posterior distribution over the latent parameter given the data is expressed as
@@
The likelihood term $p(\mathcal{D}\mid\lambda)$ is constructed from two components: i) the data likelihood $p(y_{i}\mid t_{i},x_{i},\lambda)$ captures the relationship between the inputs $t_{i},x_{i}$ and outputs $y_{i}$. ii) and the likelihood of the predicted solution satisfying the governing equations. The latter is calculated based on the residuals of eq. 1 when the model predicts the parameter and the solution. We refer to this residual as $\mathcal{R}_{\text{PDE}}$ hereafter.
@@
where $\sigma_{R}^{2}$ is the variance of the residual noise. $\mathcal{R}_{\text{PDE}}$ can either be simplified as $\sum_{i}|F(t_{i},x_{i},y_{i},\lambda)|$ or modified to include other constraints such as the boundary and initial conditions. The corresponding log-likelihood of the residuals is
@@
We sample the posterior $p(\lambda\mid\mathcal{D})$ in equation eq. 3 using variational inference (VI). Considering an approximate posterior distribution $q(\lambda)$, we want to minimize the Kullback-Leibler (KL) divergence:
@@
Substituting eq. 3 for $p(\lambda\mid\mathcal{D})$ and rearranging terms, yields the well-known Evidence Lower Bound (ELBO) objective function:
@@
Defining the expectation on $\lambda\sim q(\lambda)$ :
[…]
and separating $p(\mathcal{D}\mid\lambda)$ using the joint likelihood, eq. 10 becomes
@@
Equation 12 can be evaluated using Monte Carlo sampling of the latent variable $\lambda$ from the posterior $q(\lambda)$. The observation and residual noise variances $\sigma_{y}^{2}$ and $\sigma_{R}^{2}$ can be considered constants or learned during training. This enables a flexible uncertainty quantification method that can be utilized to estimate model or data uncertainty. We provide examples for different uncertainty estimation strategies in the experiments section 3.
The KL divergence term in eq. 12 can be approximated empirically or analytically: If both the prior $p_{0}(\lambda)$ and the variational distribution $q(\lambda)$ are assumed to be normally distributed, with means $(\mu_{0}=0,\mu_{q})$ and variances $(\sigma_{0}^{2}=1,\sigma_{q}^{2})$, the KL divergence can be written as:
@@ 2.2 Model architecture
Figure 1: DeepBayONet feedforward architecture with trunk and branch networks. The trunk network learns the PDE parameters, while the branch network predicts the forward solution of the problem.
We used a deep operator network (DeepONet) [33, 32] to parametrize $\mathcal{M}_{\theta}(t,x;\lambda)$. This architecture consists of two components: the branch network and the trunk network. The branch network approximates the forward solution $y(t,x)$ of the PDE. The trunk network processes independent prior samples from a standard normal distribution and transports them to samples of the approximate posterior $q(\lambda)$. The trunk and branch exchange information using an element-wise product before the output layers [30]. See fig. 1 for a schematic representation of different components of the architecture. Our proposed network, referred to as DeepBayONet, uses the DeepONet architecture to estimate the forward solution and parameters of the PDE at the same time:
@@
where $W_{0}$ and $b_{0}$ are the output layer weights and bias, $\sigma(\cdot)$ is the activation function, and $\odot$ denotes the element-wise product. The branch network $\mathsf{b}(t,x)$ and trunk network $\mathsf{\tau}(\lambda)$ are general parametrized functions and can be implemented using any neural network architectures, including convolutional, recurrent, or transformer networks. The entire set of trainable parameters in the model is denoted by $\theta$.
2.3 DeepBayONet Uncertainty estimation
@@
Uncertainty estimation is critical in deep neural networks, particularly in scientific modeling and prediction tasks, to assess the reliability of predictions and the robustness of the model. Two sources of uncertainty can be considered: data uncertainty arises from inherent noise in the observations, reflecting variability that cannot be reduced even with additional data. Under normal distribution assumptions, this type of uncertainty is often captured by modeling the observation error (co)variance. Specifically in deep learning, the predicted output of the network is augmented to include both the mean-prediction as well as the variance of the predicted output [34].
In contrast, model uncertainty reflects the network’s lack of flexibility or insufficient training around the underlying data distribution and can be reduced by incorporating additional data or improving the model architecture. Bayesian neural networks (BNNs) and dropout are among various methods that can address epistemic uncertainty [34] by learning a distribution over the model parameters, allowing the model to express uncertainty in regions of the input space where the model’s predictions are ambiguous. These two sources of uncertainty provide complementary information: data uncertainty quantifies noise in the data, while model uncertainty assesses the model’s confidence in its predictions.
@@
In DeepBayONet we have the ability to quantify both sources of uncertainty. Data uncertainty can be captured through a learnable observation noise variance $\sigma_{y}^{2}(x;\theta)$ which appears in the data likelihood term of the ELBO loss eq. 12 and accounts for the inherent variability in the measurements. By treating $\sigma_{y}^{2}$ as a learnable parameter, the model can adaptively estimate the data uncertainty based on the inputs. High values of $\sigma_{y}^{2}$ indicate regions with significant observation noise, leading to reduced confidence in the predictions.
Model uncertainty in DeepBayONet is inferred by randomly perturbing the latent representation $h(t,x;\lambda)=b(t,x;\lambda)$ of the forward solution. This latent representation is taken from the deep layers in the branch network before the output layer (see fig. 1). This mechanism of perturbation bears a conceptual resemblance to dropout [35], widely used for regularization as well as epistemic (model) uncertainty quantification. Dropout introduces noise during training by randomly setting activations to zero with a probability sampled from a Bernoulli distribution [35]. Let ${h}\in\mathbb{R}^{d}$ denote the hidden activations of a layer. During dropout, each element $h_{i}$ is multiplied by a binary random variable $z_{i}$ drawn independently from a Bernoulli distribution:
$z_{i}$ $\sim\mathrm{Bernoulli}(p),\quad\text{for }i=1,\dots,d,$ (15a)
$\tilde{{h}}$ $={z}\odot{h},$ (15b)
[…]
Our framework adopts a similar stochastic perturbation principle to that of dropout, but with a key difference: DeepBayONet introduces continuous perturbations to the solution representation. These perturbations are drawn from a distribution over perturbations learned by the trunk network:
$\lambda$ $\sim\mathcal{N}(0,\mathbf{I}),$ (16a)
$\tilde{{h}}$ $=\tau(\lambda)\odot{h}.$ (16b)
This continuous perturbation mechanism has two significant implications. First, it ensures that the model’s predictions generalize well by preventing the branch network from overfitting to deterministic solutions. Second, it provides a means to quantify the discovered parameters’ uncertainty, as the variability in $\mathsf{\tau}(\lambda)$ represents the model’s confidence in its parameter estimates. The posterior distribution $q(\lambda)$ over the latent parameters can be empirically estimated using the trunk network to understand the model’s confidence in the predicted parameter.
Finally, we note that if one is interested in estimating the residual noise variance $\sigma_{R}^{2}$, associated with the residuals of the governing equation eq. 8 It can also be learned as a parametric function of the network.
The DeepBayONet architecture naturally integrates regularization and uncertainty quantification in a unified framework, leveraging the interaction between the branch and trunk networks to achieve both objectives.
@@ 2.4 Practical Aspects of the Loss Function
It is possible to numerically optimize the cost function eq. 12 through Monte Carlo sampling over $t,x,\lambda$. When input samples are drawn from the training data, the data likelihood term will guide the model to predict the known forward solution. When inputs are not in the dataset, the residual likelihood trains the model to adhere to physical constraints and the PDE equations. Simultaneously, the entropy term in eq. 12 encourages the posterior of the learned parameter to remain normally distributed.
A common practice in the PINNs literature is to scale different components of the loss function to give importance to certain aspects and de-emphasize others. We follow the same practical technique and assign different weights to the terms in eq. 12.
Assuming $\mathcal{M}_{\theta}(t_{i},x_{i};\lambda)=\tilde{y}_{i}$ is the predicted PDE solution, the total loss is expressed as:
$\mathcal{L}=w_{\text{Interior}}L_{\text{Interior}}+w_{\text{IC}}L_{\text{IC}}+w_{\text{BC}}L_{\text{BC}}+w_{\text{STD}}L_{\text{STD}}+w_{\text{Data}}L_{\text{Data}},$ (17)
where
@@
$L_{\text{Data}}(\theta,\lambda;X_{\text{Data}})$ $=\frac{1}{|X_{\text{Data}}|}\sum_{(x)\in X_{\text{Data}}}\frac{\left\|\mathcal{M}_{\theta}(t,x;\lambda)-y(t,x)\right\|^{2}}{\sigma_{y}^{2}},$ (18d)
$L_{\text{STD}}(\theta,\lambda)$ $=\left\|\log{\sigma_{y}^{2}}\right\|,$ (18e)
where we have expanded the log-likelihood of the data term in eq. 12 to also include any boundary and initial conditions we may want to enforce for the learned solution. [^1] The loss component in eq. 18e was found empirically to be simpler than evaluating the exact KL-divergence. The intuition behind it is that the log of the variance of the predicted solution needs to be of moderate magnitude. Too large and too small variances are discouraged.
The weights $w_{\text{Interior}},w_{\text{IC}}$, etc., are inevitably hyperparameters that balance the influence of the respective loss components, and their choice can highly affect the quality of the learned solution. Adaptively choosing these weights is an active area of research in PINNs [36, 37]. In this paper, we have used a combination of grid search and hand-tuning of hyperparameters.